OpenStudy (anonymous):

Find the LCM of the largest 2-digit integer, the square of eleven and the square of nine.

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@jim_thompson5910 ,@Compassionate ,@tkhunny

ganeshie8 (ganeshie8):

whats the largest 2-digit integer ?

OpenStudy (anonymous):

99*99=9801

OpenStudy (anonymous):

squARE Of 11 is132 and 9 is 81

ganeshie8 (ganeshie8):

thats right ! i would have worked it by prime factorization..

OpenStudy (anonymous):

lol square of 11 is 121

OpenStudy (anonymous):

ya @Yahoo!

OpenStudy (anonymous):

so answer is 9801

ganeshie8 (ganeshie8):

99 = 9.11 = 3^2.11 11^2 = 11^2 9^2 = 3^4

OpenStudy (anonymous):

121*81=9801

ganeshie8 (ganeshie8):

LCM = 3^4*11^2 = 9801

OpenStudy (anonymous):

How many numbers divisible by each of the numbers 21, 36 and 66 are there, such that they are less than 10, 000?

ganeshie8 (ganeshie8):

the least number thats divisible by given set of numbers is their LCM

OpenStudy (anonymous):

2772

ganeshie8 (ganeshie8):

thats the LCM

ganeshie8 (ganeshie8):

so any number of form 2772k will divisible all 3 numbers 21,36 and 66

ganeshie8 (ganeshie8):

how many numbers are there of that form under 10,000 ?

OpenStudy (anonymous):

9999/2772 =remainder 1683 9999-1683=ans?

ganeshie8 (ganeshie8):

nope, you need to find the multiples of 2772 : 2772x1 = 2772 2772x2 = ? 2772x3 = ? 2772x4 = ?

ganeshie8 (ganeshie8):

stop when you exceed 10,000

OpenStudy (anonymous):

its upto 3

OpenStudy (anonymous):

8316

ganeshie8 (ganeshie8):

yes, there are exactly 3 numbers less than 10,000 that will be divisible by all 21,36 and 66